OpenStudy (anonymous):

What is an equation of a parabola with the given vertex (-2,5) and passes through these points: (-3,13) and (-1, 13)

OpenStudy (zale101):

This is the standard vertex formula y=a(x-h)^2+k where the vertex is h,k

OpenStudy (anonymous):

-2,5

OpenStudy (zale101):

that means h=-2 and k=5

OpenStudy (anonymous):

yes

OpenStudy (zale101):

y=a(x-(-2)^2+5

OpenStudy (anonymous):

i know but what is a

OpenStudy (anonymous):

a=?

OpenStudy (zale101):

substitute (x,y) = (-3,13) 13=a(3+2)^2+5 13 = a(5)² + 5 13 = a(25) + 5 13 = 25a + 5 8=25a a=8/25